# Data Assetization via Resources-decoupled Federated Learning

### v1

@@ 5. Conclusion and Future works
In this work, we explore a method for data assetization through FL. We first propose the resource-decoupled FL framework to address the challenges of data resource heterogeneity and the mismatch between data and computing resources. Building on this, to address the potential issue of participant willingness to engage, we formulate a Tripartite Stackelberg model. Our theoretical analysis guarantees the existence of a Stackelberg-Nash equilibrium that optimizes global utility and achieves the goal of maximizing data assets. We use the contribution of data to the model as a standard for assessing data quality, incentivizing data owners to enhance their contributions to model quality in order to achieve better personal gains. Our comprehensive experiments show that our method can achieve optimal global utility in various scenarios. In future work, we aim to expand the experimental scope to validate the effectiveness of our algorithm across more datasets and further optimize the global model by incorporating a selection mechanism for the model owner.
@@ A.1. Proof of Lemma 2
In order to give the Nash equilibrium optimal solution for n-bit data owners, we give the following proof. According to Equ.13, we set this first derivative equal to 0 for each data owner, then we get:
(20) $\frac{\eta}{\sum^{N}_{i=0}q_{i}}-\frac{\eta q_{n}}{(\sum^{N}_{i=0}q_{i})^{2}}=\frac{\lambda\rho}{f_{n}}$
@@
(23) $\sum^{N}_{i=0}q_{i}=\frac{(N-1)\eta}{\sum^{N}_{i=1}\frac{\lambda\rho}{f_{i}}}$
@@
Finally, we can get the optimal strategy for the data owner as shown in Equ.15.
A.2. Attached Experimental Diagram
For the specific selected data owners, each of one’s utility $U_{n}$ increases with the report deviation of $f_{n}$ on MNIST and CIFAR-10 datasets, as shown Figs. 7a and 7b.

### v2

@@ 5. Conclusions and Future works
In this work, we explore a method for data assetization through FL. We first propose the resource-decoupled FL framework to address the challenges of data resource heterogeneity and the mismatch between data and computing resources. Building on this, to address the potential issue of participant willingness to engage, we formulate a Tripartite Stackelberg model. Our theoretical analysis guarantees the existence of a Stackelberg-Nash equilibrium that optimizes global utility and achieves the goal of maximizing the value of data assets. We use the contribution of data to the model as a standard for assessing data quality, incentivizing data owners to enhance their contributions to model quality in order to achieve better personal gains. Our comprehensive experiments show that our method can achieve optimal global utility in various scenarios. In future work, we aim to expand the experimental scope to validate the effectiveness of our algorithm across more datasets and further optimize the global model by incorporating a selection mechanism for the model owner.
@@ A.1. Proof of Lemma 2
In order to give the optimal Nash equilibrium solution for $N$ data owners, we give the following proof. According to Equ.13, we set this first derivative equal to 0 for each data owner, then we get:
(20) $\frac{\eta}{\sum^{N}_{i=1}q_{i}}-\frac{\eta q_{n}}{(\sum^{N}_{i=1}q_{i})^{2}}=\frac{\lambda\rho}{f_{n}}$
@@
(23) $\sum^{N}_{i=1}q_{i}=\frac{(N-1)\eta}{\sum^{N}_{i=1}\frac{\lambda\rho}{f_{i}}}$
@@
Thus, the utility function is quasi-concave, so a unique Nash equilibrium exists.
A.2. Proof of Theorem 1
We can follow the proof of Theorem 1.
The existence and uniqueness of SNE can be deduced by the property that the strategy space is a convex and compact subspace of the Euclidean space while the utility functions are concave. In the case of the model owner, it can be justified that the optimal utility $U_{s}$ is obtained only at $\eta^{*}$ derived by direct derivation due to the strictly concave property of Equ. 7, leading to the first inequality in Def. 2 holding uniquely at $\eta^{*}$, according to Lemma 5. Similarly, a unique Nash equilibrium can be justified, and the second inequality holds for every data owner at $q_{n}^{*}$, according to Lemma 4. However, for privacy reasons, we don’t split the data in the computing center; instead, we use the Gale-Shapley algorithm to achieve one-time matching with the data owner according to the optimal $d_{m}^{*}$, according to Lemma 3. According to (Teo et al., 2001; Gokhale et al., 2024), the Gale-Shapley algorithm satisfies the relaxed Nash equilibrium, thus satisfying the third inequality. Therefore, SNE exists since the set of inequalities in Equ. 9 can be satisfied at $\left\langle\eta^{*},Q^{*},G^{*}\right\rangle$.
A.3. Gale-Shapley Algorithm
In this work, we adopt the Gale-Shapley algorithm to find the match between the computing centers and the data owners. The details and pseudocode are shown in Alg.3.
Algorithm 3 Gale-Shapley
0:  $D=\{D_{1},D_{2},...,D_{N}\}$, $C=\{C_{1},C_{2},...,C_{M}\}$, and $\sigma_{m}$
0:  $G=\{G_{1},G_{2},...,G_{N}\}$
1:  Gets $C$ sorted in descending order by $\sigma_{m}$ represented as $C^{*}$
2:  for each $D_{n},n\in[1,N]$ do
3:     $proposal=C^{*}$
4:  end for
5:  for each $C_{m},m\in[1,M]$ do
6:     Calculate $d_{m}^{*}$ according to Equ.11
7:     for each $D_{n},n\in[1,N]$ do
8:        $preference=|d_{m}^{*}-x_{n}^{*}|$
9:     end for
10:     sort $preference$ in ascending order
11:  end for
12:  while $\exists D_{n},n\in[1,N]$ $s.t.$ $G_{n}=\varnothing$  do
13:     for $D_{n},n\in[1,N]$ $s.t.$ $G_{n}=\varnothing$ do
14:        $c\leftarrow$ $\max(proposal)$ to whom $D_{n}$ has not proposed yet
15:        if $\exists D_{j},j\in[1,N]$ $s.t.$ $G_{j}=c$ then
16:           if  $preference[c]$ of $D_{n}$ is greater $D_{j}$ then
17:              $G_{n}\leftarrow c$, $G_{j}\leftarrow\varnothing$
18:           end if
19:        else
20:           $G_{n}\leftarrow c$
21:        end if
22:     end for
23:  end while
24:  return  $G$
A.4. Experiments Setting
Our training utilizes a widely used and well-established model architecture: a Convolutional Neural Network (CNN). For the simpler MNIST dataset, we use CNN, which consists of two convolutional layers and two fully connected layers. For datasets CIFAR-10 and CIFAR-100, we use CNN composed of three convolutional layers and four fully connected layers. The FL training process involves 100 iterations, including three local epochs of client training. Cross-entropy loss is employed for all classification tasks. Meanwhile, we adopt the Adam optimizer with a learning rate of 0.001. All experiments are conducted on PyTorch 1.11.0 and RTX 4090. In the absence of special instructions, the specific parameters of the experiment in this paper are set in Table 3.
Table 3. Experiment Setup
Parameters Values
$\lambda$ (Market regulating factor) 1
$\rho$ (Unit training payment of $D$) 1
$f_{n}$ (Data quality of training by $D_{n}$) $U\sim[0,1]$
$\varepsilon$ (Unit training cost of $C$) 1
$\sigma_{m}$ (Computational power of $C_{m}$) $U\sim[0,1]$
$\alpha$ (Adjustment factor of model quality) 5
$f_{m}$ (Data quality of $C_{m}$ calculate during training) $U\sim[0,1]$
$\beta$ (Learning rate for local model parameter update) $51e-$ 5
A.5. Supplementary Results of (RQ2)
In MNIST and CIFAR-10, we randomly selected three data owners from $N,N=10$ for the experiment. For each selected data owner, we fix the values of the other data owners, varying their $f_{n}$ from $-60\%$ to $+60\%$ of the actual value of $f_{n}$, and observe the variation of $U_{n}$. The results in Figs. LABEL:sub@fig7a and LABEL:sub@fig7b show that, as the deviation ratio of data owner reported $f_{n}$ increases, $U_{n}$ increases on different datasets. However, false data quality increases the personal utility value, affecting other high-quality data owners’ utility in making a proper decision, thus affecting the final global utility.
[…]
@@ A.5. Supplementary Results of (RQ2)
Figure 8. Effect of reported $f_{n}$ to $U_{s}$ and the accuracy of global model on CIFAR-10, (a) $U_{s}$ with different proportions of data owners and different report deviation ratio $f_{n}$, and (b) the accuracy of global model with different $f_{n}$.
A.6. Experiment on Computing Center
To observe the effect of the matching and parameter analysis of computing centers, we set $M=10$ and $N=10$ and observe the matching results on the different datasets. As shown in Tables 4, 5 and 6, for each computing center, they first calculate the $d^{*}_{m}$ according to Equ.12. Then, the value of $preference$ makes the highest priority to the data owner who provides $x^{*}_{n}$ closer to $d^{*}_{m}$. At the same time, every data owner tends to choose the computing center with high $\sigma_{m}$. Here, we randomly set $\sigma_{m}\in[0,1]$ of the computing centers with o.1 step size. Experimental results show that our algorithm can obtain the approximate optimal amount of data based on $d^{*}_{m}$ and $x^{*}_{n}$. In fact, without considering the privacy issue, we can make the data center undertakes exactly $d^{*}_{m}$ of the data to optimal its utility, but this violates the privacy constraint. It should be noted that since the data owner will choose inversely according to the value of $\sigma_{m}$, the matching result is the optimal solution that satisfies the wishes of both parties. For example, in Table 4, $d_{m}$ of $C_{1}$ is 1327, and $x_{n}$ of $D_{1}0$ is also 1327, but they do not match since $\sigma_{1}$ is only 0.1.
Table 4. Matching results, $U_{m}$, and related parameters on MNIST
$D_{1}$ $D_{2}$ $D_{3}$ $D_{4}$ $D_{5}$ $D_{6}$ $D_{7}$ $D_{8}$ $D_{9}$ $D_{10}$
$C$ $C_{5}$ $C_{1}$ $C_{2}$ $C_{9}$ $C_{8}$ $C_{6}$ $C_{4}$ $C_{7}$ $C_{10}$ $C_{3}$
$\sigma_{m}$ 0.5 0.1 0.2 0.9 0.8 0.6 0.4 0.7 1.0 0.3
$d_{m}$ 798 1327 1195 269 401 666 930 533 136 1062
$x_{n}$ 584 446 577 157 445 719 1120 556 136 1327
$U_{m}$ 0.069 0.127 0.110 0.015 0.008 0.021 0.091 0.040 0.000 0.088
Table 5. Matching results, $U_{m}$, and related parameters on CIFAR-10
$D_{1}$ $D_{2}$ $D_{3}$ $D_{4}$ $D_{5}$ $D_{6}$ $D_{7}$ $D_{8}$ $D_{9}$ $D_{10}$
$C$ $C_{5}$ $C_{7}$ $C_{6}$ $C_{9}$ $C_{8}$ $C_{4}$ $C_{10}$ $C_{1}$ $C_{3}$ $C_{2}$
$\sigma_{m}$ 0.5 0.7 0.6 0.9 0.8 0.4 1.0 0.1 0.3 0.2
$d_{m}$ 1183 756 969 329 542 1396 115 2037 1610 1823
$x_{n}$ 1795 820 199 408 715 2036 115 148 183 179
$U_{m}$ 0.078 0.040 0.052 0.013 0.023 0.093 0.000 0.043 0.042 0.093
Table 6. Matching results, $U_{m}$, and related parameters on CIFAR-100
$D_{1}$ $D_{2}$ $D_{3}$ $D_{4}$ $D_{5}$ $D_{6}$ $D_{7}$ $D_{8}$ $D_{9}$ $D_{10}$
$C$ $C_{7}$ $C_{5}$ $C_{8}$ $C_{2}$ $C_{6}$ $C_{9}$ $C_{10}$ $C_{3}$ $C_{1}$ $C_{4}$
$\sigma_{m}$ 0.7 0.5 0.8 0.2 0.6 0.9 1.0 0.3 0.1 0.4
$d_{m}$ 726 1152 513 1791 939 300 87 1578 2004 1365
$x_{n}$ 714 1807 452 110 440 273 86 148 98 2003
$U_{m}$ 0.026 0.081 0.033 0.048 0.032 0.015 0.000 0.112 0.096 0.076